Can anyone help with this trig HW?


In Exercises 39-46, find the unit vector that has the same direction as the unit vector v.

39. V= 6i

40. v = 5j

41. v=3i-4j

42. v= 8i-6j

43. v = 3i -2j

44. v = 4i-2j

45. v = i + j

46. v = i - j


Are the Vectors Equal?

1. Find ║u║ 

2. Find ║v║

3. Is u = v?

Explain

Q1. The points on u are (-1,2) and (4,6)

The points on v are (0,0) and (5,4)


Q2. The points on u are (-4,6) and (0,0)

The points on v are (-2,5) and (2,-1)


Q3. The points on u are (-1,1) and (5,1) 

The points on v are (-2,-1) and (4,-1)


Q4. The points on u are (-3,3)and (-3,-2) 

The points on v are (3,1) and (3,-4)


1
Expert's answer
2021-05-17T03:50:03-0400

"1.\ V = 6i\\
|V| = \sqrt{6²} = 6\\
u = 6i/6 = i"


"2.\ V= 5j\\
|V| = \sqrt{5²} = 5
\\ u = 5j/5 = j"


"3.\ V = 3i-4j\\
|V| = \sqrt{3²+4²} = 5"


"u = \dfrac{3i-4j}{5} = \dfrac{3i}{5} -\dfrac{4j}{5}"


"4. \ V = 8i-6j\\
|V| = \sqrt{8²+6²} = 10"


"u = \dfrac{8i-6j}{5} = \dfrac{4i}{5} -\dfrac{3j}{5}"


"5.\ V = 3i-2j\\
|V| = \sqrt{3²+2²} = \sqrt{13}"


"u = \dfrac{3i}{\sqrt{13}}- \dfrac{2j}{\sqrt{13}}"


"6.\ V = 4i-2j\\
|V| = \sqrt{4²+2²} = 2\sqrt5"


"u =\dfrac{2i}{√5} -\dfrac{j}{\sqrt{5}}"


"7.\ V = i+j\\
|V| = \sqrt{1²+1²} = \sqrt2"


"u =\dfrac{i}{\sqrt{2}}+\dfrac{j}{\sqrt{2}}"


"8.\ V = i-j\\
|V| = \sqrt{1²+1²} = \sqrt2"


"u =\dfrac{i}{\sqrt{2}} -\dfrac{j}{\sqrt{2}}"


Q1.

1. "║u║ = \sqrt{41}"

2. "║u║ = \sqrt{41}"

3. Yes


Q2.

1. "║u║ =2\sqrt{13}"

2. "\u2551v\u2551= 2\u221a13"

3. Yes


Q3.

1. "\u2551u\u2551 = 6"

2. "\u2551v\u2551= 6"

3. Yes


Q4.

1. "\u2551u\u2551 = 5"

2. "\u2551v\u2551= 5"

3. Yes